OpenStudy (anonymous):

. 17 – 6 × 10 ¸ 2 + 12

OpenStudy (anonymous):

Follow the order of operations PEMDAS

OpenStudy (anonymous):

17-6=11x10=110+12 =122

OpenStudy (anonymous):

^ DID not follow the order of operations!!!

OpenStudy (anonymous):

So @OutSpoken IS WRONG!!!!!

OpenStudy (anonymous):

posib ans A. 27.8 B. 59 C. 67 D. –1

OpenStudy (anonymous):

You need to rewrite your question because it's very confusing !

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

well first off rewrite the questions because its confusing

OpenStudy (anonymous):

i did follow the order of operations i just missed the 2 and the question is confusing